Suppose the price of bagels in Allentown is currently $1.30 per bagel. There are 10 low-cost bakeries and 10 high-cost bakeries that can produce bagels, each of which has the supply function

  

low-cost bakery: Qslow-cost= 200P - 180
high-cost bakery: Qshigh-cost = 200P - 280

          (These individual supply functions apply in the short run and the long run.)

a. Which bakeries will be active when the price is $1.30?

     (Click to select)Only high-cost bakeries will be active both high- and low-cost bakeries will be actively low-cost bakeries will be active.

b. If the price rises to $1.80, what will be the market supply in the short run?

     .

     What will be the market supply in the long run?
Answer #1

a) low-cost bakeries will be active. This is because the quantity is positive at this price (200*1.3 - 180 = 80 units but 200*1.3 - 280 = -20 units)

b) In the short run, continuing with the fact that new firms cannot enter, only low-cost bakeries will be active. Hence market supply will be 200*1.80 - 180 = 180*10 = 1800 units in the short run.

At P = 1.80, Qslow-cost = 200*1.8 - 180 = 180 units, Qshigh-cost = 200*1.8 - 280 = 80 units so total market supplied in the long run = 180*10 + 80*10 = 2600 units.

Hence, in the short run market supply is 1800 units but in the long run it is 2600 units.